# How Students Use AI Feedback Matters: Experimental Evidence on Physics Achievement and Autonomy

### v1

@@ 1 Introduction
Regarding the learning outcome, both learning ability and academic achievement are crucial, representing related but distinct constructs. Academic achievement denotes the level of mastery demonstrated by assessments, whereas learning ability refers to the combination of cognitive, metacognitive, and motivational capacities that enable individuals to acquire and apply knowledge effectively across various contexts (UNESCO, 2022). Within this framework, autonomy plays a pivotal role. Holec (1981) defines it as "the ability to take charge of one’s own learning" (p. 3). When learners perceive autonomy in their educational environment, they are more likely to exhibit intrinsic motivation, leading to deeper engagement and improved learning outcomes (Deci & Ryan, 2000). Autonomous learners tend to actively engage in self-regulated learning (Oguguo et al., 2023), a process widely recognized as a key driver of academic success and lifelong development (Boyraz et al., 2016). These contemporary understanding of autonomy aligns with the philosophical perspective wherein moral agency arises from individuals acting according to principles they have rationally chosen for themselves (Kant, 1992).
@@ 2.1 Educational personalized recommendation
Regarding non-academic outcomes, a few studies have examined the effects on psychology, cognition, and behavior. Hsu et al. (2013) found that recommending reading materials aligned with students’ proficiency significantly reduced cognitive load. Drissi et al. (2024) combined gamification elements with recommendation features reported increased motivation. However, it lacked a distinct control condition to identify the independent effect of recommendation. For self-regulated learning, Du et al. (2024) recruited 81 graduate students from an online course and employed a three-cycle design-based research approach to develop a recommender system for promoting SRL. Through iterative refinement based on interview feedback and pre-post behavioral data, they found that the system significantly improved SRL skills but had no significant impact on academic performance. However, although the ERS developed in this study demonstrated a positive impact on learners’ self-regulation skills, its pedagogical approach notably differs from other EPR systems: it focused on recommending self-regulatory behaviors (such as goal setting or reflection strategies) rather than specific learning materials. Therefore, the findings cannot directly address questions concerning the effects of traditional content-based recommendation strategies on learning processes and outcomes. The question of how high school students respond when recommended with personalized educational content remains insufficiently addressed. In the era when GAI significantly enhances the capabilities of EPR, it is especially important to examine its comprehensive effects on learners’ academic achievement and autonomy, given their rapid growth and increasing practical influence.
@@ 2.2 Educational on-demand help
On-demand Help (ODH) refers to assistance provided in response to a learner’s proactive request for support, commonly observed in both human teaching interactions and intelligent educational systems (ref). From a Vygotskian perspective, ODH occurs when learners encounter cognitive or practical problems they cannot independently solve, prompting them to seek support from external collaborators (ref). Although seeking help and providing ODH have been longstanding educational phenomena, traditional classroom practices and educational psychology research have mostly focused on teaching methods and classroom activities, paying little attention to how students seek and receive help on an individual level. (Aleven et al., 2003). Some theories have treated "help seeking" as a distinct psychological phenomenon within education, including it as a significant variable within self-regulated learning and learner autonomy frameworks (Zimmerman, 2011). Intelligent tutoring systems and digital learning platforms, offer highly controlled environments that enable detailed observation and collection of learners’ behavioral and psychological data. This capability facilitates in-depth exploration of help-related behaviors and the underlying learning mechanisms (Yang and Stefaniak, 2023).
@@ 2.3 Learner autonomy and self-regulated learning
Learner autonomy (LA) refers to learners’ ability to take charge of and manage their own learning processes (Holec, 1981; Little, 1991), involving full responsibility for setting goals, selecting materials, deciding activities and strategies, monitoring progress, and evaluating outcomes. Autonomous learners proactively determine their learning content and methods, take complete responsibility for the learning processes and outcomes, seek assistance without complete dependency, and adjust strategies according to learning needs and contexts. Recent theoretical developments divide LA into technological, psychological, political, and social dimensions (Bensen, 2001). Among these, the technological and psychological dimensions align closely with the concept of self-regulated learning (SRL), as both describe learner control and management over the learning process, including motivation, goal setting, execution monitoring, and reflective evaluation. The political and social dimensions of LA emphasize learners’ autonomous beliefs in relationships and responsibilities, incorporating learning environments, teacher-student relationships, and broader political contexts within education and society. Specifically, the political dimension addresses learners’ perceptions of freedom, rights in educational decision-making, and their relationships with educational authorities. The social dimension examines students’ tendency to form interdependent learning relationships with those around them, including their inclination to communicate with, seek support from, and compare themselves to their teachers and classmates (Murase, 2015).
@@
Autonomous learners actively manage their learning through self-regulated strategies. Self-regulated learning is a crucial process that enables learners to actively set goals, regulate their cognitive, motivational, and behavioral strategies, and adapt to different learning environments (Zimmerman, 2011). Models including Zimmerman’s cyclical model, Pintrich’s self-regulation framework, and Winne and Hadwin’s information processing model, emphasize the dynamic nature of SRL, highlighting its role in goal setting, self-monitoring, and feedback-based adjustments (Zimmerman, 2011; Pintrich, 2000; Winne & Hadwin, 1998). Despite minor variations in perspectives, these models share a common focus on the interplay between cognition, motivation, and behavioral control as key components of SRL. Self-regulation is regarded as a crucial ability fostering both short-term academic success and long-term developmental achievements (McClelland & Cameron, 2012). In the field of STEM learning, students who develop SRL skills perform better on immediate tasks, such as problem-solving and computational thinking, and retain these skills over time, improving their ability to adapt and succeed in complex academic tasks (Tsai et al., 2013). Self-regulation interventions have also been shown to exert lasting influence beyond academic settings, with significant effects on individuals’ lifelong health and social outcomes (Pandey et al., 2018).
@@
Recent empirical studies have illuminated the relationship between students’ academic achievement levels and their SRL strategies and autonomy, suggesting differentiated instructional approaches may be beneficial. Cheng et al. (2024) conducted a study involving 66 secondary school students and 59 university students, utilizing trace data to analyze SRL processes during multi-source writing tasks. Findings indicated that lower-performing secondary students predominantly engaged in basic SRL processes like orientation, relying heavily on task instructions and rubrics, whereas higher-performing students employed more advanced strategies such as re-reading and organization, highlighting a need for scaffolding to enhance SRL skills among lower achievers. Similarly, Admiraal et al. (2024) implemented a one-year field experiment with 432 secondary students across three schools to examine the effects of autonomy-support interventions on SRL strategies. The intervention granting the highest level of student control resulted in significant improvements in task orientation, planning, and process evaluation strategies, albeit with a slight decrease in perceived autonomy support, suggesting that structured autonomy support can effectively bolster SRL among students. In higher education, Oguguo et al. (2023) surveyed 256 undergraduates to assess the impact of SRL, autonomy, and agency on academic achievement in an online research methods course. The study revealed a strong positive correlation (r = 0.862) between these variables and academic performance, with SRL, autonomy, and agency collectively accounting for 74.4% of the variance in students’ achievement, underscoring the importance of fostering these skills for academic success. Collectively, these studies suggest that lower-achieving students may benefit from more structured and supportive learning environments to develop SRL and autonomy, while higher-achieving students are better equipped to thrive in settings that offer greater autonomy. Yet, the literatures have not conclusively addressed which modes of using GAI-powered educational resources best accommodate learners at different achievement and autonomy levels, highlighting an essential area requiring further empirical clarification.
@@ 3.1.2 Procedures
Before and after the intervention, researchers administered self-regulated learning questionnaires. Pre- and post-intervention test scores were used as outcome indicators of academic achievement in subsequent analyses. The procedure is displayed in Figure 1.
@@ 3.1.3 Instrument design
In terms of emotional support, the study observed that typical workbook answers and explanations were mostly neutral and objective, whereas generative AI could offer feedback with greater emotional resonance through carefully crafted instructions and example-based prompting. Providing emotionally supportive feedback has been shown to facilitate learners’ cognitive and emotional development, including using non-judgmental language when pointing out mistakes, affirming students’ existing efforts and understanding, encouraging further attempts, and promoting an accepting attitude toward errors (Rogers et al., 2011). These principles were explicitly integrated into the workflow’s prompts, and chain-of-thought examples (Wei et al., 2022) were provided for agent training and reference.
@@ 3.1.4 Measurements
Experiment 1 used a 15-item self-regulated learning questionnaire developed by Dowson and McInerney (2004) as instrument for autonomy. The reliability of the instrument was confirmed, with Cronbach’s alpha coefficients of 0.790 for the pretest and 0.823 for the post-test.
@@ 3.1.5 Data analysis
where $Outcome_{ij}$ is the standardized exam grade or self-regulated learning questionnaire result of student $i$ in class $j$; $TreatmentA_{ij}$ and $TreatmentB_{ij}$ are binary variables indicating treatment assignments for student $i$ in class $j$; $Pretestscore_{ij}$ controls for student performance on physics, capturing the exam score in the previous term of student $i$ in class $j$; $PreSRL_{ij}$ controls for the self-regulated learning self-report questionnaire score gathered before the experiment of student $i$ in class $j$; and Class $j$ controls class $j$’s class level fixed effect. Robust standard errors are calculated at the individual level, which is the unit of randomization.
@@ 3.2 Experiment 2: autonomous usage of GAI-powered on-demand help
In the semester following Experiment 1, Experiment 2 was conducted within the same grade at the same school to address research questions 2 and 3. Experiment 2 included two experimental groups—on-demand help with full student control and on-demand help with student-system shared control–and a non-intervention control group to examine how fully learner-controlled on-demand help and algorithm-learner jointly controlled help-seeking, under AI-generated content scenarios, influence students’ academic performance and self-regulated learning.
@@ 3.2.1 Participants and sampling
In Experiment 2, participant recruitment expanded to include all students in the grade level enrolled in physics courses (n = 373), following the same informed consent and ethical principles as in Experiment 1. Among the recruited students, 124 were assigned to the on-demand help group where students have full control over feedback type selection (Group D), 124 to the shared-control help group where student can raise questions but the system decides which types of feedback are provided (Group E), and 125 to the control group (Group C2). Participants were individually randomized within each of the seven administrative classes, ensuring equal distribution among the three experimental conditions within each class. Ultimately, 266 students (Group D: n = 87; Group E: n = 88; Group C2: n = 91) provided valid pre- and post-intervention test scores and completed the pre- and post-self-regulated learning questionnaires.
@@ 3.2.5 Data analysis
where $Outcome_{ij}$ is the standardized exam grade or self-regulated learning questionnaire result of student $i$ in class $j$; $TreatmentD_{ij}$ and $TreatmentE_{ij}$ are binary variables indicating treatment assignments for student $i$ in class $j$; $Pretestscore_{ij}$ controls for student performance on physics, capturing exam score in the previous term of student $i$ in class $j$; $PreLA_{ij}$ controls for the self-regulated learning self-report questionnaire score gathered before the experiment of student $i$ in class $j$; and $Class_{j}$ controls class $j$’s class level fixed effect. Robust standard errors are calculated at the individual level, which is the unit of randomization.
@@ 4.1.1 Sample
Among all participants, 121 (Group A: n = 42; Group B: n = 38; Group C1: n = 41) submitted valid pre- and post-test scores and completed the self-regulated learning questionnaires. Students were grouped into top-, medium-, and bottom-third cohorts according to their pre-test performance. Group A contained 15 top students, Group B contained 12 top students, and Group C1 contained 14 top students. Fourteen medium-third students were randomly distributed to Group A, 10 to Group B, and 17 to Group C1. Thirteen bottom-third students were distributed to Group A, 16 to Group B, and 10 to Group C1.
Across the entire sample, one-way ANOVA results revealed no statistically significant differences among Groups A, B, and C1 in terms of pretest academic performance, F(2, 118) = 0.09, p = .914, or pretest self-regulated learning, F(2, 118) = 1.93, p = .150. Consistent patterns were observed within each performance-level subsample. Specifically, no significant group differences were found in pretest academic performance within the top-third subsample, F(2, 38) = 1.67, p = .202; the medium-third subsample, F(2, 38) = 0.81, p = .453; or the bottom-third subsample, F(2, 36) = 0.74, p = .484. Likewise, no significant differences were observed in pretest self-regulated learning within the top-third subsample, F(2, 38) = 1.90, p = .163; the medium-third subsample, F(2, 38) = 0.08, p = .928; or the bottom-third subsample, F(2, 36) = 0.71, p = .498. These findings suggest that the three groups were comparable at baseline across both academic and self-regulatory domains.
The assumption of homogeneity of variance for both pretest academic performance and pretest self-regulated learning was met across all groups and performance-level sub-samples, as indicated by non-significant Bartlett’s tests (p > .05 for all comparisons). These results, combined with the non-significant ANOVA findings, indicate that the three experimental groups were well-matched at baseline, demonstrating strong comparability and equivalence both in the overall sample and within each sub-sample prior to the intervention.

### v2

@@ 1 Introduction
Regarding the learning outcome, both learning ability and academic achievement are crucial, representing related but distinct constructs. Academic achievement denotes the level of mastery demonstrated by assessments, whereas learning ability refers to the combination of cognitive, metacognitive, and motivational capacities that enable individuals to acquire and apply knowledge effectively across various contexts (UNESCO, 2022). Within this framework, autonomy plays a pivotal role. Holec (1981) defines it as "the ability to take charge of one’s own learning" (p. 3). When learners perceive autonomy in their educational environment, they are more likely to exhibit intrinsic motivation, leading to deeper engagement and improved learning outcomes (Deci & Ryan, 2000). Autonomous learners tend to actively engage in self-regulated learning (SRL) (Oguguo et al., 2023), a process widely recognized as a key driver of academic success and lifelong development (Boyraz et al., 2016). These contemporary understanding of autonomy aligns with the philosophical perspective wherein moral agency arises from individuals acting according to principles they have rationally chosen for themselves (Kant, 1992).
@@ 2.1 Educational personalized recommendation
Regarding non-academic outcomes, a few studies have examined the effects on psychology, cognition, and behavior. Hsu et al. (2013) found that recommending reading materials aligned with students’ proficiency significantly reduced cognitive load. Drissi et al. (2024) combined gamification elements with recommendation features reported increased motivation. However, it lacked a distinct control condition to identify the independent effect of recommendation. For SRL, Du et al. (2024) recruited 81 graduate students from an online course and employed a three-cycle design-based research approach to develop a recommender system for promoting SRL. Through iterative refinement based on interview feedback and pre-post behavioral data, they found that the system significantly improved SRL skills but had no significant impact on academic performance. However, although the ERS developed in this study demonstrated a positive impact on learners’ self-regulation skills, its pedagogical approach notably differs from other EPR systems: it focused on recommending self-regulatory behaviors (such as goal setting or reflection strategies) rather than specific learning materials. Therefore, the findings cannot directly address questions concerning the effects of traditional content-based recommendation strategies on learning processes and outcomes. The question of how high school students respond when recommended with personalized educational content remains insufficiently addressed. In the era when GAI significantly enhances the capabilities of EPR, it is especially important to examine its comprehensive effects on learners’ academic achievement and autonomy, given their rapid growth and increasing practical influence.
@@ 2.2 Educational on-demand help
On-demand Help (ODH) refers to assistance provided in response to a learner’s proactive request for support, commonly observed in both human teaching interactions and intelligent educational systems (ref). From a Vygotskian perspective, ODH occurs when learners encounter cognitive or practical problems they cannot independently solve, prompting them to seek support from external collaborators (ref). Although seeking help and providing ODH have been longstanding educational phenomena, traditional classroom practices and educational psychology research have mostly focused on teaching methods and classroom activities, paying little attention to how students seek and receive help on an individual level. (Aleven et al., 2003). Some theories have treated "help seeking" as a distinct psychological phenomenon within education, including it as a significant variable within SRL and learner autonomy frameworks (Zimmerman, 2011). Intelligent tutoring systems and digital learning platforms, offer highly controlled environments that enable detailed observation and collection of learners’ behavioral and psychological data. This capability facilitates in-depth exploration of help-related behaviors and the underlying learning mechanisms (Yang and Stefaniak, 2023).
@@ 2.3 Learner autonomy and self-regulated learning
Learner autonomy (LA) refers to learners’ ability to take charge of and manage their own learning processes (Holec, 1981; Little, 1991), involving full responsibility for setting goals, selecting materials, deciding activities and strategies, monitoring progress, and evaluating outcomes. Autonomous learners proactively determine their learning content and methods, take complete responsibility for the learning processes and outcomes, seek assistance without complete dependency, and adjust strategies according to learning needs and contexts. Recent theoretical developments divide LA into technological, psychological, political, and social dimensions (Bensen, 2001). Among these, the technological and psychological dimensions align closely with the concept of SRL, as both describe learner control and management over the learning process, including motivation, goal setting, execution monitoring, and reflective evaluation. The political and social dimensions of LA emphasize learners’ autonomous beliefs in relationships and responsibilities, incorporating learning environments, teacher-student relationships, and broader political contexts within education and society. Specifically, the political dimension addresses learners’ perceptions of freedom, rights in educational decision-making, and their relationships with educational authorities. The social dimension examines students’ tendency to form interdependent learning relationships with those around them, including their inclination to communicate with, seek support from, and compare themselves to their teachers and classmates (Murase, 2015).
@@
Autonomous learners actively manage their learning through self-regulated strategies. SRL is a crucial process that enables learners to actively set goals, regulate their cognitive, motivational, and behavioral strategies, and adapt to different learning environments (Zimmerman, 2011). Models including Zimmerman’s cyclical model, Pintrich’s self-regulation framework, and Winne and Hadwin’s information processing model, emphasize the dynamic nature of SRL, highlighting its role in goal setting, self-monitoring, and feedback-based adjustments (Zimmerman, 2011; Pintrich, 2000; Winne & Hadwin, 1998). Despite minor variations in perspectives, these models share a common focus on the interplay between cognition, motivation, and behavioral control as key components of SRL. Self-regulation is regarded as a crucial ability fostering both short-term academic success and long-term developmental achievements (McClelland & Cameron, 2012). In the field of STEM learning, students who develop SRL skills perform better on immediate tasks, such as problem-solving and computational thinking, and retain these skills over time, improving their ability to adapt and succeed in complex academic tasks (Tsai et al., 2013). Self-regulation interventions have also been shown to exert lasting influence beyond academic settings, with significant effects on individuals’ lifelong health and social outcomes (Pandey et al., 2018).
@@
Recent empirical studies have illuminated the relationship between students’ academic achievement levels and their SRL strategies and autonomy, suggesting differentiated instructional approaches may be beneficial. Cheng et al. (2024) conducted a study involving 66 secondary school students and 59 university students, utilizing trace data to analyze SRL processes during multi-source writing tasks. Findings indicated that lower-performing secondary students predominantly engaged in basic SRL processes like orientation, relying heavily on task instructions and rubrics, whereas higher-performing students employed more advanced strategies such as re-reading and organization, highlighting a need for scaffolding to enhance SRL skills among lower achievers. In higher education, Oguguo et al. (2023) surveyed 256 undergraduates to assess the impact of SRL, autonomy, and agency on academic achievement in an online research methods course. The study revealed a strong positive correlation (r = 0.862) between these variables and academic performance, with SRL, autonomy, and agency collectively accounting for 74.4% of the variance in students’ achievement, underscoring the importance of fostering these skills for academic success. Regarding the use of GAI, in a randomized controlled trial, Li et al. (2025) developed and tested GAI-powered SRL hints—adaptive prompts generated by GAI based on real-time analytics of learners’ SRL processes. Similar to Du et al. (2024), the design stressed on the metacognition aspect of learning instead of the cognitive (i.e. knowledge comprehension) one. The study found that participants receiving these personalized SRL hints exhibited significantly more metacognitive learning behaviors compared to control groups. Collectively, these studies suggests that lower-achieving students might benefit from more structured learning support to develop SRL and autonomy, and the latter skills can be fostered through hints and guidance generated by GAI. However, existing intervention studies on GAI-powered feedback tend to focus exclusively on SRL or autonomy, often without reporting corresponding changes in achievement. Conversely, studies that report achievement gains typically do not directly measure SRL or autonomy, instead inferring their impact through indirect reasoning. This disconnect highlights the need for systematic investigations into the interplay between GAI usage patterns, achievement and autonomy outcomes.
@@ 3.1.2 Procedures
Before and after the intervention, researchers administered SRL questionnaires. Pre- and post-intervention test scores were used as outcome indicators of academic achievement in subsequent analyses. The procedure is displayed in Figure 1.
@@ 3.1.3 Instrument design
In terms of emotional support, the study observed that typical workbook answers and explanations were mostly neutral and objective, whereas GAI could offer feedback with greater emotional resonance through carefully crafted instructions and example-based prompting. Providing emotionally supportive feedback has been shown to facilitate learners’ cognitive and emotional development, including using non-judgmental language when pointing out mistakes, affirming students’ existing efforts and understanding, encouraging further attempts, and promoting an accepting attitude toward errors (Rogers et al., 2011). These principles were explicitly integrated into the workflow’s prompts, and chain-of-thought examples (Wei et al., 2022) were provided for agent training and reference.
@@ 3.1.4 Measurements
Experiment 1 used a 15-item SRL questionnaire developed by Dowson and McInerney (2004) as instrument for autonomy. The reliability of the instrument was confirmed, with Cronbach’s alpha coefficients of 0.790 for the pretest and 0.823 for the post-test.
@@ 3.1.5 Data analysis
where $Outcome_{ij}$ is the standardized exam grade or SRL questionnaire result of student $i$ in class $j$; $TreatmentA_{ij}$ and $TreatmentB_{ij}$ are binary variables indicating treatment assignments for student $i$ in class $j$; $Pretestscore_{ij}$ controls for student performance on physics, capturing the exam score in the previous term of student $i$ in class $j$; $PreSRL_{ij}$ controls for the SRL self-report questionnaire score gathered before the experiment of student $i$ in class $j$; and Class $j$ controls class $j$’s class level fixed effect. Robust standard errors are calculated at the individual level, which is the unit of randomization.
@@ 3.2 Experiment 2: autonomous usage of GAI-powered on-demand help
In the semester following Experiment 1, Experiment 2 was conducted within the same grade at the same school to address research questions 2 and 3. Experiment 2 included two experimental groups—on-demand help with full student control and on-demand help with student-system shared control–and a non-intervention control group to examine how fully learner-controlled on-demand help and algorithm-learner jointly controlled help-seeking, under AI-generated content scenarios, influence students’ academic performance and SRL.
@@ 3.2.1 Participants and sampling
In Experiment 2, participant recruitment expanded to include all students in the grade level enrolled in physics courses (n = 373), following the same informed consent and ethical principles as in Experiment 1. Among the recruited students, 124 were assigned to the on-demand help group where students have full control over feedback type selection (Group D), 124 to the shared-control help group where student can raise questions but the system decides which types of feedback are provided (Group E), and 125 to the control group (Group C2). Participants were individually randomized within each of the seven administrative classes, ensuring equal distribution among the three experimental conditions within each class. Ultimately, 266 students (Group D: n = 87; Group E: n = 88; Group C2: n = 91) provided valid pre- and post-intervention test scores and completed the pre- and post-SRL questionnaires.
@@ 3.2.5 Data analysis
where $Outcome_{ij}$ is the standardized exam grade or SRL questionnaire result of student $i$ in class $j$; $TreatmentD_{ij}$ and $TreatmentE_{ij}$ are binary variables indicating treatment assignments for student $i$ in class $j$; $Pretestscore_{ij}$ controls for student performance on physics, capturing exam score in the previous term of student $i$ in class $j$; $PreLA_{ij}$ controls for the SRL self-report questionnaire score gathered before the experiment of student $i$ in class $j$; and $Class_{j}$ controls class $j$’s class level fixed effect. Robust standard errors are calculated at the individual level, which is the unit of randomization.
@@ 4.1.1 Sample
Among all participants, 121 (Group A: n = 42; Group B: n = 38; Group C1: n = 41) submitted valid pre- and post-test scores and completed the SRL questionnaires. Students were grouped into top-, medium-, and bottom-third cohorts according to their pre-test performance. Group A contained 15 top students, Group B contained 12 top students, and Group C1 contained 14 top students. Fourteen medium-third students were randomly distributed to Group A, 10 to Group B, and 17 to Group C1. Thirteen bottom-third students were distributed to Group A, 16 to Group B, and 10 to Group C1.
Across the entire sample, one-way ANOVA results revealed no statistically significant differences among Groups A, B, and C1 in terms of pretest academic performance, F(2, 118) = 0.09, p = .914, or pretest SRL, F(2, 118) = 1.93, p = .150. Consistent patterns were observed within each performance-level subsample. Specifically, no significant group differences were found in pretest academic performance within the top-third subsample, F(2, 38) = 1.67, p = .202; the medium-third subsample, F(2, 38) = 0.81, p = .453; or the bottom-third subsample, F(2, 36) = 0.74, p = .484. Likewise, no significant differences were observed in pretest SRL within the top-third subsample, F(2, 38) = 1.90, p = .163; the medium-third subsample, F(2, 38) = 0.08, p = .928; or the bottom-third subsample, F(2, 36) = 0.71, p = .498. These findings suggest that the three groups were comparable at baseline across both academic and self-regulatory domains.
The assumption of homogeneity of variance for both pretest academic performance and pretest SRL was met across all groups and performance-level sub-samples, as indicated by non-significant Bartlett’s tests (p > .05 for all comparisons). These results, combined with the non-significant ANOVA findings, indicate that the three experimental groups were well-matched at baseline, demonstrating strong comparability and equivalence both in the overall sample and within each sub-sample prior to the intervention.